OpenStudy (anonymous):

x^2+6-7 solve the trinomial

OpenStudy (cwrw238):

should that be 6x?

OpenStudy (anonymous):

Yes, sorry x^2+6x-7

OpenStudy (anonymous):

I would recommend using the quadratic formula.

OpenStudy (anonymous):

You would get x=-7 and x=1

OpenStudy (cwrw238):

you need 2 numbers whose product is -7 and sum +6 factors will be of the form (a +/- a)(x +/- b)

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

the numbers are +7 and -1

OpenStudy (anonymous):

@cwrw238 please double check your answers.

OpenStudy (cwrw238):

x^2 + 6x - 7 = ( x+7)(x - 1)

OpenStudy (anonymous):

@cwrw238 That's correct so far, but when you set (x+7)=0 and solve it, you get x=-7 and x=1

OpenStudy (cwrw238):

yes - i was only dealing with the factoring when i said +7 and -1 i meant the 2 numbers which go into the parentheses solutions are -7 and 1

OpenStudy (cwrw238):

- these are the solution to x^2 + 6x - 7 = 0 of course

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

Yep

OpenStudy (cwrw238):

yw